OpenStudy (nuccioreggie):

The price of an all-access ticket to the fair is $42.99 before tax. Roland bought the all-access ticket for 40% off. He then paid 7% sales tax on the discounted price. Part A: How much did Roland pay in sales tax? Show all work and steps in your solution. (5 points) Part B: What is the total amount that Roland paid for the ticket? Show all work and steps in your solution. (5 points)

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

@Z4K4R1Y4

OpenStudy (anonymous):

he payed 1.80 in tax

OpenStudy (michele_laino):

please, show to @nuccioreggie your reasoning @wardb

OpenStudy (nuccioreggie):

Please

OpenStudy (anonymous):

ok will the discounted price comes out to 25.974

OpenStudy (michele_laino):

the price payed by Roland is: \[42.99 - \left( {0.4 \times 42.99} \right) = ...?\]

OpenStudy (anonymous):

25.794 times 7%= $1.80558

OpenStudy (nuccioreggie):

25.794

OpenStudy (anonymous):

part b: he payed a total of $27.59958

OpenStudy (anonymous):

because 25.794+1.80558=$27.59958

OpenStudy (anonymous):

does that make sense?

OpenStudy (nuccioreggie):

@Michele_Laino can you explain part A annd Part B please

OpenStudy (nuccioreggie):

Is he correct for part B @Michele_Laino

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (is3535):

B